Crews worked through the weekend maintaining clean-up efforts. The prolonged rains had a positive flushing effect on the creek. Some upstream boons are collecting little to no emulsified oil. Cleanup personnel did not observe any pockets of trapped product throughout the day yesterday (Monday). Sheens have become light to moderate dull gray, with little of the rainbow sheens that were observed last week. The odor has also subsided noticeably. The NJ DEP has remained on location throughout the weekend to the present.

 Although there are reports of oiled beavers, with at least one deceased and another captured for treatment, there are no reports of any waterfowl in distress or oiled.

There remains no end date for the remediation effort and work crews are scheduled to remain on location. Residents are urged to stay out of the water and keep their pets away from the water until further notice. Any questions, please contact OEM at This email address is being protected from spambots. You need JavaScript enabled to view it..

Please click here to view the statement from DEP